# Enabling Large-Scale and High-Precision Fluid Simulations on Near-Term Quantum Computers

### v2

@@ 5 Conclusion and Outlook
In this paper, we propose a novel approach to Computational Fluid Dynamics (CFD) leveraging the capabilities of near-term quantum computers, along with several enhancements to existing methods to enable effective CFD simulations. First, we introduce a high-precision quantum linear solver suitable for Noisy Intermediate-Scale Quantum (NISQ) devices, called Iterative-QLS. This solver effectively mitigates errors that typically arise in quantum computing processes. Our simulation of a 2D Poiseuille flow on a near-term quantum computer demonstrates a relative error of less than $0.2\%$, even lower than the gate error of the quantum computer. To address the challenges in quantum state tomography, we develop a sparse tomography algorithm, which removes the dependency on quantum random access memory inherent in $l_{\infty}$-norm tomography. Second, we introduce the subspace method to scale the Iterative-QLS method for larger CFD problems, enhancing solution convergence and resulting in a quantum-classical hybrid approach. By employing the subspace method, we successfully simulate unsteady acoustic wave propagation on a $41\times 41$ grid using a NISQ computer, marking a significant advancement toward scalable quantum-enhanced CFD solutions.
@@ Appendix A Quantum hardware specification
Fig. 6 shows the chip-topology of Wukong, and the chosen qubits used in experiments of this work. The parameters of used qubits are shown in Tab. 2. The fidelity of CZ gates are shown in Tab. 3.
@@ B.1 Full algorithm of the Iterative-QLS
Algorithm 4 Iterative-VQLS
1:Linear equations $A\boldsymbol{x}=\boldsymbol{b},\boldsymbol{b}\neq\boldsymbol{0}$. Difference step size $\Delta$. Convergence monitor $\epsilon_{c}$. Convergence monitor of loss value $\epsilon_{l}$.Learning rate $\lambda$.
@@
16:     Obtain the gradient $\boldsymbol{g}_{k}=(L_{k}^{*}-L_{k})/\Delta$.
@@ B.2 Introduction of VQLS
Here, $\lambda$ is the learning rate, and $\boldsymbol{g}=\partial L/\partial\boldsymbol{\theta}$ is the gradient of the loss function. There are multiple ways to obtain the gradient. For example, one can directly observe every component of $\boldsymbol{g}$ [24] using Hadamard-Test circuit. If a small enough loss value has been found, the information of the corresponding quantum state can be extracted through quantum tomography in the end.
@@ B.3 Classical computation for loss value and its gradient
$g_{i}\approx\frac{L(\theta_{i}+\frac{\pi}{2})-L(\theta_{i}-\frac{\pi}{2})}{2}.$ (19)
@@ C.1 Full algorithm of the subspace method with Iterative-QLS
1:Linear equations $A\boldsymbol{x}=\boldsymbol{b}$. Initial value $\tilde{\boldsymbol{x}}\neq A^{-1}\boldsymbol{b}$. Precision constraint $\epsilon$. Subspace dimension $m$.
2:Calculate residual $\boldsymbol{r}=\vec{b}-A\tilde{\boldsymbol{x}}$ and $\beta=\|\boldsymbol{r}\|_{2}$.
3:if  $\beta\leq\epsilon$ then
4:     Goto line 37.
5:end if
6:$\boldsymbol{v}_{1}=\boldsymbol{r}_{0}/\beta$, $\boldsymbol{\xi}=\beta\boldsymbol{e}_{1}$.
7:for  $j\in[1,m]$ do
8:     $\boldsymbol{w}_{j}=A\boldsymbol{v}_{j}$.
9:     for  $i\in[1,j]$ do
10:         $h_{ij}=(\boldsymbol{w}_{j},\boldsymbol{v}_{i})$.
11:         $\boldsymbol{w}_{j}=\boldsymbol{w}_{j}-h_{ij}\boldsymbol{v}_{i}$.
12:     end for
13:     $h_{j+1,j}=\|\boldsymbol{w}_{j}\|_{2}$.
14:     for  $i\in[1,j-1]$ do
15:         $\begin{bmatrix}h_{i,j}\\ h_{i+1,j}\end{bmatrix}=\begin{bmatrix}c_{i}&s_{i}\\ -s_{i}&c_{i}\end{bmatrix}\begin{bmatrix}h_{i,j}\\ h_{i+1,j}\end{bmatrix}$.
16:     end for
17:     if  $h_{j+1,j}=0$ then
18:         $m=j$.
19:         break
20:     end if
21:     $\boldsymbol{v}_{j+1}=\boldsymbol{w}_{j}/h_{j+1,j}$.
22:     if  $|h_{j,j}|>|h_{j+1,j}|$ then
23:         $c_{j}=1/\sqrt{1+\tau^{2}},s_{j}=c_{j}\tau,\tau=h_{j+1,j}/h_{j,j}$.
24:     else
25:         $s_{j}=1/\sqrt{1+\tau^{2}},c_{j}=s_{j}\tau,\tau=h_{j,j}/h_{j+1,j}$.
26:     end if
27:     $h_{j,j}=c_{j}h_{j,j}+s_{j}h_{j+1,j}$.
28:     $h_{j+1,j}=0$.
29:     $\begin{bmatrix}\xi_{j}\\ \xi_{j+1}\end{bmatrix}=\begin{bmatrix}c_{j}&s_{j}\\ -s_{j}&c_{j}\end{bmatrix}\begin{bmatrix}\xi_{j}\\ 0\end{bmatrix}$.
30:     if  $|\xi_{j+1}|<\beta\epsilon$ then
31:         $m=j$.
32:         break
33:     end if
34:end for
35:Solve the subspace linear system $H\boldsymbol{y}=\boldsymbol{\xi}$ through Iterative-QLS. The matrix $H$ is composed of elements $h$.
36:Update $\tilde{\boldsymbol{x}}=\tilde{\boldsymbol{x}}+V\boldsymbol{y}$, where $V=[\boldsymbol{v}_{1},\cdots,\boldsymbol{v}_{j}]$.
37:Goto line 1.
38:Return $\tilde{\boldsymbol{x}}$.
@@ C.2 Convergence analysis
The starting residual of the equation is $\boldsymbol{r}_{0}=\boldsymbol{b}-A\boldsymbol{x}_{0}$. The equation will be converted to solve $A\boldsymbol{x}_{1}=\boldsymbol{r}_{0}$ and update the solution to $\boldsymbol{x}=\boldsymbol{x}_{0}+\boldsymbol{x}_{1}$. Therefore, this lemma is equivalent to showing that we can always provide a better solution within the $m$-th order Krylov subspace compared to $\vec{0}$ as the initial guess for the converted equation. Note that $\lVert\boldsymbol{x}-\hat{\boldsymbol{x}}\rVert\leq\mathcal{O}(e^{-(m/\kappa)^{2}})$, the Chebyshev polynomial of the first kind has already provided such a solution with better precision, given any $m$ and $b$.
@@
The sbuspace method produces a solution with $\epsilon$ error as long as QLS finds a solution better than $\boldsymbol{0}$ as the initial guess. That is, we can solve the linear system with precision $\max(\mathcal{O}(e^{-m^{2}/\kappa^{2}}),\epsilon_{q})<1$ such that $C=\langle\psi|A^{\dagger}(I-|b\rangle\langle b|)A|\psi\rangle\leq\epsilon_{q}$. The number of iterations is bounded by
@@ Appendix D Theory details of sparse tomography
For an $N$-dimension real-valued quantum state $|x\rangle=\sum_{i=0}^{N-1}x_{i}|i\rangle$, we produce a classical vector $\hat{\boldsymbol{x}}$ that satisfies $\|\boldsymbol{x}-\hat{\boldsymbol{x}}\|_{\infty}<\epsilon$ with a number of copies of the quantum state. The classical vector has $k$ nonzero entries such that $k\ll N$. Here, real-valued means that every $x_{i}$ is a real number.
@@
Given access of unitary $U$ such that $U|0\rangle=|x\rangle$ and its controlled version, there exists an algorithm that produces an $l_{\infty}$ sampling for a real-valued quantum state with $\mathcal{O}(\log d/\epsilon^{2})$ query complexity. The algorithm requires a constant number of queries to QRAM and an extra qubit.
@@
1:Error $\delta$, access to unitary $U:|0\rangle\to|x\rangle=\sum_{i\in[d]x_{i}|i\rangle}$, the controlled version of $U$, QRAM access.
2:Classical vector $\tilde{X}\in\mathbb{R}^{d}$, such that $\lVert\tilde{X}\rVert=1$ and $\lVert\tilde{X}-\boldsymbol{x}\rVert_{\infty}<\delta$.
3:Measure $N=36\ln d/\delta^{2}$ copies of $|x\rangle$ in the standard basis and count $n_{i}$, the number of times the outcome $i$ is observed. Store $\sqrt{p_{i}}=\sqrt{n_{i}/N}$ in QRAM data structure.
4:Create $N=36\ln d/\delta^{2}$ copies of the state $\frac{1}{\sqrt{2}}|0\rangle\sum_{i\in[d]}x_{i}|i\rangle+\frac{1}{\sqrt{2}}|1\rangle\sum_{i\in[d]}\sqrt{p_{i}}|i\rangle$.
@@
$|\phi\rangle=\frac{1}{2}\sum_{i\in[d]}\left(\left(x_{i}+\sqrt{p_{i}}\right)|0,i\rangle+\left(x_{i}-\sqrt{p_{i}}\right)|1,i\rangle\right)$
6:Measure both registers of each copy in the standard basis, and count $n(0,i)$ the number of time the outcome $(0,i)$ is observed.
7:Set $\sigma(i)=+1$ if $n(0,i)>0.4Np_{i}$ and $\sigma(i)=-1$ otherwise.
8:Output the unit vector $\tilde{X}$ such that $\forall i\in[N],\tilde{X}_{i}=\sigma_{i}\sqrt{p_{i}}$.
@@
Similar to $l_{\infty}$ tomography, the first step is to sample $36\log d/\epsilon^{2}$ copies of the quantum state to produce a positive valued vector to the quantum state. Note that $k<36\log d/\epsilon^{2}$ is always satisfied so that the classical vector is sparse when $36\log d/\epsilon^{2}<d$. Now the major target is to determine the relative phase of these sparse entries of the quantum state. The technique is to apply shadow tomography to compute the sign of each term.
@@
Given that each edge is successfully measured with a constant success rate $p$, we can traverse this graph randomly in $v$ steps and determine all signs successfully with a probability of $p^{v}$. Repeating this random traversal $k$ times will result in a success rate of approximately $1-\left(1-p^{v}\right)^{k}\sim 1-\left(v\left(1-p\right)\right)^{k}$. Therefore, we can choose $k=(\log\delta)/\left(\log{v\left(1-p\right)}\right)$ to ensure the algorithm succeeds with a probability of $1-\delta$, with the number of measurements being $\mathcal{O}(Mvk)$, where $M$ still represents the number of measurements for each Hamiltonian, and $\mathcal{O}(vk)$ denotes the number of Hamiltonians that need to be computed.
@@
$\mathcal{O}(Mvk)$ $=\mathcal{O}\left(\frac{18\log^{2}d\cdot(\log(36)+\log(\log d)-2\log\epsilon)}{\epsilon^{2}(\log(\log d)+12)}\right)$ (27)
$=\widetilde{\mathcal{O}}\left(\frac{\log^{2}d\cdot\left|\log\epsilon\right|}{\epsilon^{2}}\right).$
This result implies that the algorithm should sample $\widetilde{\mathcal{O}}({\epsilon^{-2}}{\log^{2}d})$ times to generate all required routes in this fully connected graph.
@@
$\mathcal{O}(M\log(vk))$ $=\mathcal{O}\left(\frac{\log d\cdot\log\left(\frac{\log d}{\epsilon^{2}}\right)}{\epsilon^{2}}\right)$ (28)
$=\widetilde{\mathcal{O}}\left(\frac{\log d}{\epsilon^{2}}\right).$
@@
This complexity aligns with the complexity in the first step of the algorithm. Therefore, the sparse-tomography method can be executed with an explicit algorithm, with time complexity also being $\mathcal{O}(\log d/\epsilon^{2})$.
Appendix E Additional experiments on the scalability of Iterative-QLS
In this section, we will analyze the scalability of Iterative-QLS which acts as a fundamental tool in the subspace method. A major concern for adopting a real NISQ quantum computer is whether a noisy solution is sufficient to produce high-precision results. As analyzed in C.2, the convergence of the subspace method is guaranteed when each inner solver provides a rough solution, namely $\epsilon_{q}<1$. Therefore, in this section, we will empirically show that the Iterative-QLS can be used to produce the result at a certain amount of error.
The performance of Iterative-QLS to solve 2D incompressible Poiseuille flow running on a noise-free quantum circuit simulator is studied first, which indicates the scalability under ideal conditions. Fig. 8(a) and (b) respectively show the variation of residuals and loss values. It can be seen that iterative VQLS can successfully solve the system with 2 qubits but fails when using more than 2 qubits, even though a noticeable decrease in loss values can still be observed.
Noticing that as the number of qubits increases, the condition number of linear systems also grows. We utilize Iterative-QLS to solve the corresponding preprocessed linear systems. After preprocessing, Iterative-QLS can always successfully solve the system, and more significant decreases in loss values can be observed in some training steps as shown in Fig. 8(c) and (d).
(a)(b)(c)(d)
Figure 8: Results of iterative VQLS running on virtual machines. (a) Residual change with increasing iteration steps and (b) loss value change with training steps while ones solve the linear systems directly. (c) Residuals change with increasing iteration steps and (d) change of loss value with training steps while ones solve the linear systems after preprocessing.
It can be seen that in some other training steps, the loss values of 3-qubit and 4-qubit are consistently difficult to effectively reduce. After excluding the influence of noise, it can only be considered that this difficult-to-train phenomenon is caused by the limitations of expression of ansatz. Regardless, the iterative VQLS demonstrates its scalability under ideal conditions, although this scalability is highly sensitive to the condition number of the linear system, and an increase in the condition number will rapidly weaken the solving capability.
Fig. 9 illustrates the performance of Iterative-QLS in practice. Qualitatively, Fig. 9 the results on NISQ devices align with those obtained on the virtual machine. The best results are achieved with 2 qubits, and Iterative-QLS fails to correctly solve the system when using more than 2 qubits. As shown in Fig. 9(c) and (d), after preprocessing, Iterative-QLS demonstrates a converging trend in all cases. These consistent findings provide encouraging evidence that Iterative-QLS can exhibit scalability consistent with expectations under ideal conditions.
(a)(b)(c)(d)
Figure 9: Results of Iterative-QLS without precondition. (a) Residuals change with increasing iteration steps and (b) loss values change with training steps while ones solve the linear systems directly. (c) Residual’s change with increasing iteration steps and (d) loss value’s change with training steps after preprocessing.
A quantitative comparison between Fig. 8(d) and Fig. 9(d) reveals that the minimum loss value of Iterative-QLS is nearly an order of magnitude larger than the results obtained on virtual machines. This indicates a lack of precision in the solution at each iteration step of Iterative-QLS, ultimately leading to a significant decrease in convergence speed as illustrated in Fig. 8(c) and Fig. 9(c). While this difference in convergence speed is not pronounced with 2 qubits, it rapidly increases as the number of qubits grows. After 9 iterations on virtual machines, all residuals can decrease to around $10^{-5}$, whereas on the NISQ device, after 9 iterations the residual of 4-qubits can only decrease to around $3\times 10^{-3}$. This difference in convergence speed is believed to be caused by the noise in NISQ devices.
Appendix F Examples for the discussion about scalability
Here, we demonstrate two numerical examples to show how our algorithm can be enhanced by a better quantum computer. Firstly, we conduct quantum circuit simulations for Poiseuille flow on a classical computer with different qubit errors. The qubit error is defined as the strength of the depolarizing error channel appended after every quantum gate. In Fig. 10, we can see that the count of the circuit execution significantly decreases with a decreasing error rate. Therefore, our method promises a faster execution with a better qubit quality that is utilized in the experiment.
Figure 10: Circuit execution counts of different qubit errors in Iterative-QLS algorithm that reaches the same convergence criteria. Error bars are plotted using the standard error of 100 repetitions of each error configuration.
Secondly, we argue that the integration of the subspace method on the flow simulation is promising when more qubits are provided to enable a larger subspace size $m=2^{n}$ where $n$ is the qubit and $m$ is the size of the subspace. In Fig. 11, we solve a matrix of $1024$ dimension with $m$ varies from $2^{4}$ to $2^{10}$. We plot the final residual and the number of steps of subspace, which corresponds to the predicted circuit execution time on a quantum computer. The result implies the problem can be divided into two regimes. At $n\leq 8$, where the subspace size is far less than the dimension, we call it the “NISQ regime”, representing the case when we have a very limited number of high-quality qubits. At $n\geq 9$, where the subspace size approaches the matrix size, we call it the “FTQC regime”, representing the case when we are able to put a full flow field on a quantum computer.
In the NISQ regime, we find the number of subspace steps follows a nearly exponential dependence on the number of qubits, while the system converges to a similar residual. This represents that the investment of every additional qubit will result in exponential computing power. In the FTQC regime, the subspace size approaches the dimension of the original matrix. As a result, the residual decreases significantly because it almost becomes a direct inversion algorithm. The number of subspace steps also decreases significantly when it enters such a regime. This represents that even in the FTQC regime, it is meaningful to explore the potential of integration of the subspace method and quantum computing before direct inversion is allowed.
Refer to caption NISQ Regime FTQC Regime

### v3

@@ 5 Conclusion and Outlook
In this paper, we propose a novel approach to CFD leveraging the capabilities of near-term quantum computers, along with several enhancements to existing methods to enable effective CFD simulations. First, we introduce a high-precision quantum linear solver suitable for NISQ devices, called Iterative-QLS. This solver effectively mitigates errors that typically arise in quantum computing processes. Our simulation of a 2D Poiseuille flow on a near-term quantum computer demonstrates a relative error of less than $0.2\%$. To address the challenges in quantum state tomography, we develop a sparse tomography algorithm, which removes the dependency on quantum random access memory inherent in $l_{\infty}$-norm tomography. Second, we introduce the subspace method to scale the Iterative-QLS method for larger CFD problems, enhancing solution convergence and resulting in a quantum-classical hybrid approach. By employing the subspace method, we successfully simulate unsteady acoustic wave propagation on a $41\times 41$ grid using a superconducting quantum computer, marking a significant advancement toward scalable quantum-enhanced CFD solutions.
We highlight the importance of utilizing a hybrid scheme. If the entire $41\times 41$ grid is simulated entirely on a quantum computer, one should employ at least 13 high-quality qubits. However, in our experiments, two noisy qubits are sufficient to produce high-precision and large-scale results. Our method significantly reduces the cost of simulation on a quantum computer to fewer qubits and effectively suppresses the noise. Therefore, our proposed method provides a feasible way to scale the existing method to large-scale fluid simulation with near-term quantum computers, especially those with limited qubit numbers and gate performance.
The last question is whether our method is a scalable method to achieve potential quantum speedup. The subspace method, as a quantum-classical hybrid approach, allocates the subspace problem to quantum computers, leaving the remaining part computed on the classical method. By numerical examples in E, we show that the number of subspace iteration steps decreases significantly by investing every one more qubit. Then, one can expect quantum computers can provide speedup in the subspace problem at a certain amount of qubits, which will be left as future work.
@@ Appendix A Quantum hardware specification
Fig. 8 shows the chip-topology of Wukong, and the chosen qubits used in experiments of this work. The parameters of used qubits are shown in Tab. 1. The fidelity of CZ gates is shown in Tab. 2.
@@ B.1 Full algorithm of the Iterative-QLS
Algorithm 4 Iterative-QLS
1:Linear equations $A\boldsymbol{x}=\boldsymbol{b},\boldsymbol{b}\neq\boldsymbol{0}$. Difference step size $\Delta$. Convergence criteria $\epsilon_{c}$. Convergence criteria of loss value $\epsilon_{l}$.Learning rate $\lambda$.
@@
16:     Obtain the gradient $\boldsymbol{g}_{k}$ using parameter-shift rule.
@@ B.2 Introduction of VQLS
Here, $\lambda$ is the learning rate, and $\boldsymbol{g}=\partial L/\partial\boldsymbol{\theta}$ is the gradient of the loss function. There are multiple ways to obtain the gradient. For example, one can directly observe every component of $\boldsymbol{g}$ [30] using the Hadamard-Test circuit. If a small enough loss value has been found, the information of the corresponding quantum state can be extracted through quantum tomography in the end.
@@ B.3 Classical computation for loss value and its gradient
$g_{i}=\frac{L(\theta_{i}+\frac{\pi}{2})-L(\theta_{i}-\frac{\pi}{2})}{2}.$ (21)
@@ C.1 Full algorithm of the subspace method with Iterative-QLS
1:Linear equations $A\boldsymbol{x}=\boldsymbol{b}$. Convergence criteria $\epsilon$. Subspace dimension $m$.
2:Give an initial solution $\tilde{\boldsymbol{x}}$.
3:Obtain the residual $\boldsymbol{r}=\boldsymbol{b}-A\tilde{\boldsymbol{x}}$ and $\beta=\|\boldsymbol{r}\|_{2}$.
4:if  $\beta\leq\epsilon$ then
5:     Goto line 37.
6:end if
7:$\boldsymbol{v}_{1}=\boldsymbol{r}_{0}/\beta$, $\boldsymbol{\xi}=\beta\boldsymbol{e}_{1}$.
8:for  $j\in[1,m]$ do
9:     $\boldsymbol{w}_{j}=A\boldsymbol{v}_{j}$.
10:     for  $i\in[1,j]$ do
11:         $h_{ij}=(\boldsymbol{w}_{j},\boldsymbol{v}_{i})$.
12:         $\boldsymbol{w}_{j}=\boldsymbol{w}_{j}-h_{ij}\boldsymbol{v}_{i}$.
13:     end for
14:     $h_{j+1,j}=\|\boldsymbol{w}_{j}\|_{2}$.
15:     for  $i\in[1,j-1]$ do
16:         $\begin{bmatrix}h_{i,j}\\ h_{i+1,j}\end{bmatrix}=\begin{bmatrix}c_{i}&s_{i}\\ -s_{i}&c_{i}\end{bmatrix}\begin{bmatrix}h_{i,j}\\ h_{i+1,j}\end{bmatrix}$.
17:     end for
18:     if  $h_{j+1,j}=0$ then
19:         $m=j$.
20:         break
21:     end if
22:     $\boldsymbol{v}_{j+1}=\boldsymbol{w}_{j}/h_{j+1,j}$.
23:     if  $|h_{j,j}|>|h_{j+1,j}|$ then
24:         $c_{j}=1/\sqrt{1+\tau^{2}},s_{j}=c_{j}\tau,\tau=h_{j+1,j}/h_{j,j}$.
25:     else
26:         $s_{j}=1/\sqrt{1+\tau^{2}},c_{j}=s_{j}\tau,\tau=h_{j,j}/h_{j+1,j}$.
27:     end if
28:     $h_{j,j}=c_{j}h_{j,j}+s_{j}h_{j+1,j}$.
29:     $h_{j+1,j}=0$.
30:     $\begin{bmatrix}\xi_{j}\\ \xi_{j+1}\end{bmatrix}=\begin{bmatrix}c_{j}&s_{j}\\ -s_{j}&c_{j}\end{bmatrix}\begin{bmatrix}\xi_{j}\\ 0\end{bmatrix}$.
31:     if  $|\xi_{j+1}|<\beta\epsilon$ then
32:         $m=j$.
33:         break
34:     end if
35:end for
36:Solve the subspace linear system $H\boldsymbol{y}=\boldsymbol{\xi}$ through Iterative-QLS. The matrix $H$ is composed of elements $h$.
37:Update $\tilde{\boldsymbol{x}}=\tilde{\boldsymbol{x}}+V\boldsymbol{y}$, where $V=[\boldsymbol{v}_{1},\cdots,\boldsymbol{v}_{j}]$.
38:Goto line 1.
39:Return $\tilde{\boldsymbol{x}}$.
@@ C.2 Convergence analysis
The starting residual of the equation is $\boldsymbol{r}_{0}=\boldsymbol{b}-A\boldsymbol{x}_{0}$. The equation will be converted to solve $A\boldsymbol{x}_{1}=\boldsymbol{r}_{0}$ and update the solution to $\boldsymbol{x}=\boldsymbol{x}_{0}+\boldsymbol{x}_{1}$. Therefore, this lemma is equivalent to showing that we can always provide a better solution within the $m$-th order Krylov subspace compared to $\boldsymbol{0}$ as the initial guess for the converted equation. Note that $\lVert\boldsymbol{x}-\hat{\boldsymbol{x}}\rVert\leq\mathcal{O}(e^{-(m/\kappa)^{2}})$, the Chebyshev polynomial of the first kind has already provided such a solution with better precision, given any $m$ and $b$.
@@
The subspace method produces a solution with $\epsilon$ error as long as QLS finds a solution better than $\boldsymbol{0}$ as the initial guess. That is, we can solve the linear system with precision $\max(\mathcal{O}(e^{-m^{2}/\kappa^{2}}),\epsilon_{q})<1$ such that $C=\langle\psi|A^{\dagger}(I-|b\rangle\langle b|)A|\psi\rangle\leq\epsilon_{q}$. The number of iterations is bounded by
@@ Appendix D Theory details of sparse tomography
For a $N$-dimensional real-valued quantum state $|x\rangle=\sum_{i=0}^{N-1}x_{i}|i\rangle$, we produce a classical vector $\hat{\boldsymbol{x}}$ that satisfies $\|\boldsymbol{x}-\hat{\boldsymbol{x}}\|_{\infty}<\epsilon$ with a number of copies of the quantum state. The classical vector has $k$ nonzero entries such that $k\ll N$. Here, real-valued means that every $x_{i}$ is a real number.
@@
Given access of unitary $U$ such that $U|0\rangle=|x\rangle$ and its controlled version, there exists an algorithm that produces an $l_{\infty}$ sampling for a real-valued quantum state with $\mathcal{O}(\log N/\epsilon^{2})$ query complexity. The algorithm requires a constant number of queries to QRAM and an extra qubit.
@@
1:Error $\epsilon$, access to unitary $U:|0\rangle\to|x\rangle=\sum_{i\in[N]}x_{i}|i\rangle$, the controlled version of $U$, QRAM access.
2:Classical vector $\tilde{X}\in\mathbb{R}^{N}$, such that $\lVert\tilde{X}\rVert=1$ and $\lVert\tilde{X}-\boldsymbol{x}\rVert_{\infty}<\epsilon$.
3:Measure $M=36\log N/\epsilon^{2}$ copies of $|x\rangle$ in the standard basis and count $m_{i}$, the number of times the outcome $i$ is observed. Store $\sqrt{p_{i}}=\sqrt{m_{i}/M}$ in QRAM data structure.
4:Create $M=36\log N/\epsilon^{2}$ copies of the state $\frac{1}{\sqrt{2}}|0\rangle\sum_{i\in[N]}x_{i}|i\rangle+\frac{1}{\sqrt{2}}|1\rangle\sum_{i\in[N]}\sqrt{p_{i}}|i\rangle$.
@@
$|\phi\rangle=\frac{1}{2}\sum_{i\in[N]}\left(\left(x_{i}+\sqrt{p_{i}}\right)|0,i\rangle+\left(x_{i}-\sqrt{p_{i}}\right)|1,i\rangle\right)$
6:Measure both registers of each copy in the standard basis, and count $n(0,i)$ the number of times the outcome $(0,i)$ is observed.
7:Set $\sigma(i)=+1$ if $m(0,i)>0.4Mp_{i}$ and $\sigma(i)=-1$ otherwise.
8:Output the unit vector $\tilde{X}$ such that $\forall i\in[M],\tilde{X}_{i}=\sigma_{i}\sqrt{p_{i}}$.
@@
Similar to $l_{\infty}$ tomography, the first step is to sample $36\log N/\epsilon^{2}$ copies of the quantum state to produce a positive valued vector to the quantum state. Note that $k<36\log N/\epsilon^{2}$ is always satisfied so that the classical vector is sparse when $36\log N/\epsilon^{2}<N$. Now the major target is to determine the relative phase of these sparse entries of the quantum state. The technique is to apply shadow tomography to compute the sign of each term.
@@
Given that each edge is successfully measured with a constant success rate $p$, we can traverse this graph randomly in $v$ steps and determine all signs successfully with a probability of $p^{v}$. Repeating this random traversal $k$ times will result in a success rate of approximately $1-\left(1-p^{v}\right)^{k}\sim 1-\left(v\left(1-p\right)\right)^{k}$. Therefore, we can choose $k=(\log\epsilon)/\left(\log{v\left(1-p\right)}\right)$ to ensure the algorithm succeeds with a probability of $1-\epsilon$, with the number of measurements being $\mathcal{O}(Mvk)$, where $M$ still represents the number of measurements for each Hamiltonian, and $\mathcal{O}(vk)$ denotes the number of Hamiltonians that need to be computed.
@@
$\mathcal{O}(Mvk)$ $=\mathcal{O}\left(\frac{18\log^{2}N\cdot(\log(36)+\log(\log N)-2\log\epsilon)}{\epsilon^{2}(\log(\log N)+12)}\right)$ (29)
$=\widetilde{\mathcal{O}}\left(\frac{\log^{2}N\cdot\left|\log\epsilon\right|}{\epsilon^{2}}\right).$
This result implies that the algorithm should sample $\widetilde{\mathcal{O}}({\epsilon^{-2}}{\log^{2}N})$ times to generate all required routes in this fully connected graph.
@@
$\mathcal{O}(M\log(vk))$ $=\mathcal{O}\left(\frac{\log N\cdot\log\left(\frac{\log N}{\epsilon^{2}}\right)}{\epsilon^{2}}\right)$ (30)
$=\widetilde{\mathcal{O}}\left(\frac{\log N}{\epsilon^{2}}\right).$
@@
This complexity aligns with the complexity in the first step of the algorithm. Therefore, the sparse-tomography method can be executed with an explicit algorithm, with time complexity also being $\mathcal{O}(\log N/\epsilon^{2})$.
Appendix E Examples for the discussion about scalability of the subspace method
Here, we demonstrate examples to show the scalability of the subspace method. The integration of the subspace method on the flow simulation is promising when more qubits are provided to enable a larger subspace size $m=2^{n}$ where $m$ is the size of the subspace, and $n$ corresponds to the number of qubits utilized in the internal QLS. In Fig. 10, we solve $5043$-dimensional tridiagonal matrices, the same size as the acoustic wave propagation problem presented in the main text, with $m$ varies from $2^{4}$ to $2^{10}$. We test two matrices with different condition numbers and plot the number of subspace steps after reaching the same convergence criteria. In the context of the integration of the subspace method with a QLS, the number of steps corresponds to the number of internal QLS executions on a quantum computer.
The results show an exponential decay in subspace steps with respect to the number of qubits, which represents that the subspace method could run faster if sending a larger subspace problem into the quantum solver. Therefore, in this example, the subspace method has shown great scalability: one more qubit investment would effectively get an exponential reward in terms of speed, showing the potential ability for quantum speedup.
(a)(b)